Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Tuesday, November 24, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 24, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/24/98 (Tuesday)

Segment 2B (Castaic to Sylmar): Report for 11-24-98. Crews spent most of the day pulling wire for the vault electrical system at Romero Cyn. In addition, crews completed the forms for the future electrical box required for the vault tie-in. At Feedmill Rd, near the Santa Clara river, crews placed a 6" reamer on the directional drill to widen the bore hole. Spoils were removed from Weldon Cyn and dumped near Calgrove Blvd to act as backfill for crews working Wed. Further, at Weldon Cyn crews released compressed air from the test head after running a pig through the pipe from Castaic. Trenching crews completed 200' of ditch to EMP 72.64 while pipe was set in the ditch to near EMP 72.23. Welding and X-ray crews continued work near EMP 72.25. North of Sierra Hwy 650' of asphalt was completed to EMP 72.64 and crews began removing shoring and debris from the right of way south of Sierra Hwy.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 24 November 1998: 3A: The bore at Tuxford progressed slowly, with crews gain ~30 feet. Welding, x-ray, and coating spent the day working on the pipe for Tujunga Wash Bridge crossing. Excavation crews dug more on the north side of the wash. Backfill, padding, and fiber optic conduit crews neared completion of the south side of Van Nuys Boulevard, including the removal of shoring from the bore site. The Pacoima Wash Block Valve was placed today. Fiber optic crews continued to pull cable in Glendale. 3C: Backfill continued their work on restoring the slopes and cleaning up along the right of way north of Roxford Street. Welding, x-ray, and coating continued their work on the tie in area to the aqueduct crossing. Ditch crews completed excavating the ditch up to the block valve under Balboa overpass. Crews working on the cover for the high speed channel worked on placing steel for more concrete.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page